HUBBLE SPACE TELESCOPE DAILY REPORT #2213 PERIOD COVERED: 0000Z (UTC) 09/21/98 - 0000Z (UTC) 09/22/98 Daily Status Report as of 265/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed NIC/2 7834 (A Search for Giant Planets Around Very Young Nearby Late-type Dwarfs) The Near Infrared Camera (Camera Two) was used to make observations of the very young, late-type, dwarf star GJ182 as part of a search for large planets. Very young late-type dwarfs at distances closer than 30 pc are excellent sites to provide a direct detection of extrasolar planets. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Three Sets of WF/PC-2 7620 (WF/PC-2 CYCLE 7 Standard Darks) The WF/PC-2 was used to make a dark calibration. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. Over an extended period these data will also provide a monitor of radiation damage to the CCDs. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed STIS/CCD/MA2 7347 (Direct Imaging of Betelgeuse) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to make observations of the super giant star Betelgeuse (a.k.a., BD+75D325 or Alpha Orionis). Our UV images of Alpha Orionis obtained with the FOC in 1995 {Cycle 5} are the first direct images of the surface of a star other than the Sun. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.5 Completed STIS/CCD 7928 (STIS Spectroscopic CCD Flatfielding -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make CCD flat field calibration images in the STIS spectroscopic mode. The calibrations were completed as planned, and no problems were reported. 1.6 Completed WF/PC-2 6673 (Black Holes and Gas Disks in a Complete Sample of Radio Loud UGC Ellipticals) The WF/PC-2 was used to make observations of radio load, elliptical galaxy UGC-01004 in order to search for possible black holes and gas disks. Hydrogen-Alpha and broad band observations were taken to determine the luminosity profiles and stellar populations (on the tens of parsec scale), and detect nuclear regions or disks of gas and dust. As described in the previous report and HSTAR 6687, during this proposal, the acquisition defaulted to fine lock backup on one FGS only. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.7 Completed Three Sets of STIS/MA1 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. As described in the previous report and HSTAR 6687, during the three iterations of this proposal, the acquisition defaulted to fine lock backup on one FGS only. Otherwise, the observations were completed as planned, and no further problems were reported. 1.8 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. As described in the previous report and HSTAR 6687, during all iterations of this proposal, the acquisition defaulted to fine lock backup on one FGS only. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.9 Completed STIS/CCD/MA2 8070 (Throughput of NUV filters) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to make observations of HZ4 in order to measure the throughput of the near ultraviolet filters. The purpose is to measure the absolute sensitivity of the NUVMAMA imaging modes with the CN182, CN270, MGII, and CIII filters. A wave length calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Four Sets of STIS/CCD 8080 (STIS Spectroscopic CCD Flatfielding -- Revised) The Space Telescope Imaging Spectrograph (CCD) was used to make CCD flat fielding calibration observations in the spectroscopic mode. The observations were completed as planned, and no anomalies were reported. 1.11 Completed NIC/1/2/3 8079 (Mapping the Light Scatter in the NICMOS Coronagraphic Hole {PSF Recovery Orbits}) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of star GL905 in order to map the scattering of light within the NICMOS coronagraphic hole. The targets was dithered about the low scatter point to map out the scatter about the hole. The observations were completed as planned, and no problems were reported. 1.12 Completed NIC/1/2/3 7886 (NICMOS Snap Shot Survey of Early-Type Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the early-type galaxy NGC-524 as part of a snap-shot broad-band imaging survey. Since almost all early-type galaxies harbor dust, these images which are less sensitive to extinction and will provide superior light profiles for dynamical studies of the cores of these galaxies. The observations were completed as scheduled, and no problems were reported. 1.13 Completed Two Sets of NIC/1 7248 (Spectroscopy and Polarimetry of the Beta Pictoris) The Near Infrared Camera (Camera One) was used to make spectrographic and polarimetric observations of the star Beta Pictoris. Little is known about the Beta Pictoris disk within 50 Astronomical Units of the central star. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed Three Sets of STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. The observations were completed as scheduled, and no problems were reported. 1.15 Completed WF/PC-2/FOC/96 6576 (HST imaging of the two most distant radio galaxies) The WF/PC-2 and the Faint Object Camera (f/96) were used to make observations of distant radio galaxy Gal-0143+3. High-resolution imaging is required of this object with the HST in Lyman-Alpha and continuum. During this proposal the loss of lock documented in 2.1 and HSTAR 6689 occurred, temporarily taking down the take data flag. Otherwise, the observations were completed as planned, and no further problems were reported. 1.16 Completed Three Sets of NIC/2/3 8082 (NICMOS pure parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. During the first iteration of this proposal the loss of lock documented in 2.1 and HSTAR 6689 occurred, temporarily taking down the take data flag. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.17 Completed Six Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations take in parallel with the prime science observations. During the second iteration of this proposal the loss of lock documented in 2.1 and HSTAR 6689 occurred, temporarily taking down the take data flag. Otherwise, the observations were completed as planned, and no further problems were reported. 1.18 Completed NIC/2 7220 (Imaging of Quasar Host Galaxies) The Near Infrared Camera (Camera Two) was used to make direct imaging observations of the host galaxy for the broad absorption line quasar QSO0248-00. The observations were completed as planned, and no anomalies were reported. 1.19 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.20 Completed Nine Sets of WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.21 Completed WF/PC-2 6681 (Evolution of Pre-Main Sequence Circumstellar Nebulosity) The WF/PC-2 was used to make observations of DL-Tau to study the evolution of pre-main sequence circumstellar nebulosity. Initial HST results from imaging of disks around pre-main sequence stars provide compelling justification to observe a broader sample of young stars. The observations were completed as planned, and no anomalies were reported. 1.22 Completed NIC/1/2/3 7898 (Snapshot Survey of Proto-planetary Nebulae) The Near Infrared Camera (Cameras One, Two and Three) was used to make snapshot survey observations of the protoplanetary nebula IRAS23304. This is to help discern the structure of the nebulae and the evolutionary sequence that produces axially symmetric planetary nebulae. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 9 A loss of lock occurred at 264/155715Z, resulting in a new #44 command at the acquisition of signal. The proposals described in 1.15, 1.16, 1nd 1.17 were affected by this event. HSTAR 6690 describes the acquisition at 265/030912Z which failed when the search radius limit was exceeded on both FGSs 1 and 3. The prior full maneuver updates had succeeded, and the attitude had acceptable error vectors. Three proposals occurring during the period of the next report were affected: WF/PC-2 6827, NIC 8082, and STIS 7948. Scheduled Reacquisitions : 3 Successful : 3 2.2 FHST Updates: Scheduled : 24 Successful : 24 2.3 Operations Notes: During the day yesterday, the flight operations team moved their operations from the STOCC area to the revamped SEER area. No problems occurred during the transition. Using ROP SR-1, the SSR EDAC error counter was reset four times. Per ROP NS-3, the NSSC-1 status buffer was dumped and reset at 264/0146Z. HSTAR 6688 describes an FHST map that failed at 264/151646Z when there were no stars in tracker 2. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH